log_3 (x+1)-log_3 x= log_3 4
Answer by stanbon(75887)

log_3 (x+1)-log_3 x= log_3 4
----------
log3[(x+1)/x] = log3(4)
-----
(x+1)/x = 4
-----
x+1 = 4x
3x = 1
x = 1/3
============
Cheers,
Stan H.
============